the rental fee for a moving truck is $50 plus $2 per mile. Aida paid a total of$150. Write an equation to determine,x, the number of miles she drove..
Answer by John10(297)

The equation will be a linear one.
$2 will be the variable cost (slope)
$50 is a fixed cost which is one-time charge (y-intercept)
The linear equation has standard form: y = mx + b
y = 2x + 50. where y is the total cost ($150)
Determine the number of miles x:
2x + 50 = 150
2x = 100
x = 100/2
x = 50 miles
John10:)